Weekly LIPAc highlights (28/March-1/April)

Injector conditioning campaign

After several long runs (each one of ~10hours of CW operation) the Injector CW campaign with the 11mm PE is now considered completed. The last few days of operation were dedicated to collect data for the beam profile measurements and their analyses are on-going (see below profiles). As the RFQ conditioning campaign needs to be extended, the way forward for the injector CW campaign was agreed with CEA as follows. At first the validation of the new EMU will take place with its calibration using the 11mm PE. Then, following a short maintenance of the source (also changing the BN disk), the campaign will proceed with the 12mm PE with the goal to repeat the CW operation. The target in any case is to complete this CW campaign by mid-June, when RF-RFQ conditioning should also be completed.

Extracted current=142-145 mA, Intermediate Electrode Voltage=18 kV

Beam profile measured from solenoid magnet SOL1 @ 250A to 370A and from SOL1 @450A to 600A for 5%DC, 50%DC, and CW

RF-RFQ conditioning campaign

While the RF-RFQ conditioning reached 15%DC at nominal (for deuteron acceleration) cavity voltage of 132kV and 35%DC at 80% level cavity voltage of 106kV, it was stopped on March 30 due to vacuum interlock events. Preliminary investigation pointed out that the time needed by the system to recover the RFQ base vacuum after the vacuum peaks follows the same trend than the temperatures measured in correspondence of the RFQ couplers ceramic windows (see below graphs). Consequently, leaks at the RF couplers viton O-rings are suspected. Intensive leak checks have been performed, first by using Kr gas sprayed in pockets around the windows to be detected by the RFQ RGAs and then by using Helium, following a more standard He leak test by isolating the cryopump and using a mobile trolley. Following the increase of DC during the last months of conditioning campaign, the ceramic windows may also have experienced metallization. Further inspection of the couplers will tell us more if changing ceramic windows and O-rings should be considered. Even though spares are available on-site, the diagnosis of the issue (with the great support of INFN) is our first focus to clearly identify the coupler to be opened because the intervention will be quite invasive (with the need to disassemble the coaxial lines and disconnect them in the pit for couplers located on the north side).

Vacuum and couplers temperature trends observed during the vacuum events in the RFQ

HEBT/BD commissioning

The remote sessions with CIEMAT are progressing well. In particular the new radiation resistant RGA installed in the HEBT is now fully operative (see the below screenshot of its OPI). The noise issues at the HAL02 accelerometer in the BD was solved and the final debugging of the mass flow controllers and cRIO is on-going. The monitoring of the BPMs is ongoing after the new PV gateway service through the Data-diode was started. It appears that an overrun memory issue was causing the segmentation fault observed in the last BPMs operations. These activities will speed up when two CIEMAT colleagues together with the 3 new colleagues from Granada University will join us on-site as soon as May.

Control System

Meanwhile, the Control system supported the modification of the OPI for the isolation valve of the injector vacuum sector 2, while the modification of the chopper interlock module is on-going.

HEBT RGA operational with new OPI

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (7/March-11/March)

In the first months of 2022 the operations on-site are focused on the CW conditioning campaigns for the injector and the RFQ.

The injector conditioning campaign makes steady and quick progresses: the commissioning with 11mm plasma electrode reached 155mA of total extracted beam current in CW already several times. Fine tuning of several parameters is on-going to improve the stability of the beam (decrease the sparks, current fluctuations, improve emittance etc.) and determine a stable working point for long term CW operation.

The RF-RFQ conditioning suffered a stop in February because of a serious damage at the circulator of the RF-RFQ chain#1b. After abnormal operations and high number of arcs during conditioning in the chain#1b, the circulator was opened and the connection point between the inner plate of the circulator and the matching section of coax line was found to be broken. The EU and QST teams on-site have reacted calmly and with great spirit of collaboration and proceeded to: analyse the accident, determine the possible causes, inspect carefully all other circulators connections and define an action plan with associated risk analyses. The supplier of the circulator was contacted and we are in the process to order spare parts to repair the circulator on-site. In parallel the procurements of spares circulators will be launched. A study has been performed and associated documents and procedures have been approved to resume conditioning of the RFQ by using 7 RF-RFQ chains (out of the 8 available, i.e. with chain#1b being off). In these days a dummy load has been installed on the coaxial line of chain#1b to manage the reflected power during RF injection and RF-RFQ conditioning has just resumed.

It is difficult to have a clear idea about the impact on the schedule: the repair of the circulator on-site is planned in April and assuming that RF-RFQ conditioning will be resumed with 8 chains, we might reach CW operation by beginning of June. This will leave a couple of months for beam operation before the yearly summer maintenance break.

We had also several remote sessions with CIEMAT support to solve punctual issues (like a problem on a limit switch of a motor driven drain valve for the Beam Dump cartridge in the vault, to install the SW for a newly installed radiation resistant RGA in the HEBT) and to keep progressing in the finalization of the vacuum machine state for the HEBT.

The Diagnostic team has been working to recover full operability of the Beam Position Monitors and is currently also checking the SEM grids. The Fluorescence Profile Monitor installed in the HEBT have been moved out of the vault for a necessary recalibration that is ongoing.

We had support of a subcontractor on-site to finalize the maintenance of cooling skid (with replacement of mechanical seals in several pumps) and some re-installation work of N2 gas system in the deaerator loop of the Beam Dump skid.

Injector operation record with stable CW beam extraction

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (24/January-28/January)

The experimental campaign at the end of 2021 was successful. The first stage of Phase B+ was completed successfully by December 20 with D+ beam transported at low duty cycle and low current up to the BD. Beam operation was carried out during the day and RF-RFQ was switched to conditioning mode during the night. The duty cycle of RF-RFQ conditioning reached CW at the cavity voltage of 101 kV at the last moment of our activity on Sunday 19 December 2021.

At first, 100% was tried at 20:19 of 19 December with 20ms pulse length and 20ms repetition period (i.e. 50Hz). One event of coupler temperature over the interlock limit of 65°C at the 2A tapered cone section occurred at 20:35, however CW was kept until around 23:00.

At around 00:30 of 20 December, 100% was reached again but was not so stable.

At 07:38 of 20 December, 99.5% (mostly CW) was reached and kept it until 08:13: this was the last trial before stopping the system. See RF-RFQ conditioning summary plot for some details.

On December 20 all the systems were progressively stopped and winter maintenance with planned interventions started. We had a specialized company onsite for maintenance on the cooling skids and replacement of mechanical seals on the pumps of RF skid, RFQ skid and injector skid. After the interventions a water leak at the backup pump of injector skid could not be solved and another maintenance is planned in February. Resins for the water quality control have also been replaced in MEBT and HEBT skids. During the first week of break an alignment survey was performed in the vault for checking the difference of LEBT, RFQ and MEBT positioning with those observed in December 2020 and March 2021: the results show some systematical changes that are being analyzed carefully.

The maintenance on the injector and RF-RFQ systems was completed on 16 and 25 January respectively. The first part of 2022 will be dedicated to the injector and RF-RFQ CW campaigns. We target for the time being end of March for completing both CW campaigns.

The injector underwent standard maintenance and the Plasma Electrode was replaced with the 11mm size as agreed. The chopper was disassembled specifically for the CW campaign. The Emittance Meter Unit was checked. After the PPS test the HV conditioning started. 100 keV D+ beam extraction from the 11 mm plasma electrode is on-going for reaching 100 % DC at current higher than 150 mA: 120 mA in 10 ms pulse width over 300 ms period was reached as the first target. A trigger signal to synchronize the acquisition camera with the source pulse was implemented.

In the RF system a new SW for the LLRF autocalibration in the master chain was installed and confirmed to be operational; also the White Rabbit Grand Master was updated to suppress the desynchronization issue. In the RFQ pick-up sensors the calibration of data acquisition was completed and the data are being checked by INFN. After inspection of the DC pickup port of SSPA-A for buncher #2 (that showed signal issues last year) a test was carried out but it is confirmed that a new pick-up should be installed. The intervention is planned next week.

In the BD skid O2 control is unavailable due to N2 gas generator failure, and steady circulation through BD cartridge is to be avoided (currently the skid is kept running bypassing the cartridge). Currently we are facing a calibration issue in the pressure sensors of the BD skid and we are investigating the issue with help of CIEMAT. Remote sessions have resumed since few weeks as Japan has renewed the entry ban and CIEMAT mission is forced to be postponed.

All Fluorescence Profile Monitors (FPMs) have been disassembled and will be recalibrated (also with remote support by CIEMAT).

Beam Position Monitors (BPMs) Clock Board issue is under investigation and procurement of spares is being launched.

In the pictures below the beam operation from Local Control Room and Central Control Room

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (29/November-3/December)

These last weeks have been intense with beam operation onsite in Rokkasho.

During the daytime LIPAc beam operation was carried out with H+ beam.

During the night the RF-RFQ conditioning continued increasing the Duty Cycle. INFN joined the extra daily meeting in the evening (JST) and a plan for the night conditioning was agreed with the teams onsite.

Every Monday has been devoted to maintenance activities.

RFQ conditioning

RFQ conditioning resumed following the RFQ conditioning procedure.

Early November the RFQ conditioning reached 17.5% DC (with 3500us pulse at 50Hz). There was at this point a first series of arcs and vacuum events at the RFQ cavity, with temperature increase observed in RFQ coupler 3B.

Field flatness measurement were systematically done and INFN processed them to provide recommendations for the following steps. Upon INFN suggestion the automatic frequency following mode was implemented to continue increasing the DC.

The DC was progressively increased again to 15%  during the week from 15 to 19 November and it reached 20% DC with pulse length 4ms, repetition rate 50Hz, voltage 125kV. Over the following WE the DC was increased up to 28% (pulse length from 3ms to 5ms, voltage 130-125kV).

In the last days RFQ cavity was more unstable and reflected waves from the cavity prevented rapid increase of DC during the night.

We are aiming now to get 30%DC.

Beam operation.

Beam operation was resumed on 12 November.

Until now beam operation has been performed with H+ beam extraction. We should switch to D+ beam next week.

A number of standard beam sessions were dedicated to optimize the injector source and LEBT solenoids and steerers parameters.

Several beam sessions were dedicated to understand an unusual beam shape observed at the LEBT ACCT. The RF timing setting was scanned but was confirmed not to be the cause.

Also, the IE voltage, LEBT cone repeller voltage and H+ gas flow rate in the source were scanned, but none of them was leading to clear changes in ACCT readings.

Kr gas flow was injected at the middle of LEBT to improve the space charge compensation of H+ beam (~12mA of H+ with additional ~9mA of other molecular ions) and following scanning of the flow rate and recording of the measured pulses were done.

Beam shape investigation was also performed operating w/o chopper (limiting beam current and pulse length to protect the Dplate slits and HEBT Faraday Cup). Beam shape was confirmed as expected (rough beam w/o chopper, not square shape), but compared to the operation with chopper the beam shape observed at LEBT and MEBT ACCT is the same (i.e. beam is transported through the RFQ w/o current loss in the cavity).

No clear dependencies have been identified between the scanned parameters and the beam shape at the ACCT. Data are being processed by beam physics group.

H+ beam injection with chopper: unusual beam shape at LEBT (yellow: 20mA/V) ACCT. Part of the beam is lost in the RFQ cavity (at the MEBT ACCT observed profile is square – in blue: 15mA/V)
H+ beam injection without chopper, pulse length about 1.4 ms, repetition 1 Hz: same beam shapes at LEBT (yellow) and MEBT (blue) ACCTs (i.e. before and after the RFQ cavity)

Several sessions were dedicated to check the functionality of the beam diagnostics. The diagnostic team struggled to control the proper motion and position of the HEBT Faraday Cup (CIEMAT was providing support remotely to sort out the issue).

The debugging of the Dplate and HEBT SEM grid took some time  but was successful.

After tests with different timing settings also the FPM (fluorescence profile monitor) is now operative.

Once the diagnostic functionality was confirmed, the DPlate slit scan and the HEBT slit scan for determining the beam center and profile were systematically performed.

Experts struggled to complete the BPM (beam position monitors) calibration. Still a timing synchronization issue is under investigation but BPM are now operative.

A recabling was necessary to correct the MEBT BPM polarity. MEBT and MEL BPMs correct polarity and functionality was confirmed by kicking the beam with the steerers.

After confirmation of proper functionality of all diagnostics along the line, Beam Based Alignment was performed starting from the MEBT by steering progressively the beam and centering at the next quadrupole magnet. Some difficulties have been encountered in reproducing the beam position using the same settings of the previous day. This phenomena also needs some deeper investigations.

Last Friday we performed the centering of the beam along the line and we transported the H+ beam up to the BD: we switched then to D+ operation until the end of the operation on 17/12/2021.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Project Committee #28

21 & 22 October 2021

The 28th Project Committee (PC#28) was held at QST Rokkasho Site on October 21st and 22nd and by video conference. An update of the project status was presented as well as the outlook for 2022.

The main topics discussed were the progress since last spring’s PC#27 and the effort made in the management of the LIPAc activities, with the following highlights:

  • The beam operation in its Phase B+ configuration started after two years of installation and hardware commissioning of the equipment;
  • Remote participation access tools have been made available for LIPAc beam operation thanks to a close collaboration with the IFERC project;
  • Procurement Arrangements have been signed for the Fusion Neutron Source Design activities to be carried out between 2021 and 2025;
  • Concerns were expressed regarding the possible impact on the schedule of the entry restrictions in Japan related to Covid-19, despite an already implemented mitigation plan.
View of PC#28 with remote participants

2nd EU-JA Fusion Neutron Source workshop

04 & 05 October 2021

The 2nd Fusion Neutron Source workshop (FNS WS #2) was held remotely on October 04th and 05th.

The workshop provided a general overview of the FNS collaboration activities that are carried out in Japan and Europe within the IFMIF/EVEDA, DONES and A-FNS projects.

Additional common topics of interest have been identified. They will be further discussed in the future, in order to complement the current dense work programme already agreed among all.

Although short, the workshop aroused interest for deep technical discussions and we are looking forward to meeting in person to continue these discussions in the forthcoming workshop.

View of FNS WS #2 remote participants

Weekly LIPAc highlights (8/November-12/November)

Injector

Deuteron beam extraction for injector conditioning was carried out until 26 October and reached 30 mA @ 10% duty cycle. On 27 October proton beam extraction was started and reached quickly the current and duty level required to resume beam operation (total extracted current 20mA, duty 3%). 

RF system

One of the new Driver Anode Box (chain 1B) had an issue and was recovered by replacing it back to the old one (checks are on-going). Anode Boxes in the remaining chains will be replaced progressively in the next short term maintenances.

RF conditioning started on 28 October: in a half day, 1% duty (1ms, 10Hz) was reached. The system could very quickly ramp-up to 10% duty cycle, confirming the much improved stability of the overall RF system. Duty cycle reached 10% at 2ms, 50Hz. During last weekend at ~17% duty cycle there was an important RFQ vacuum interlock event and associated RFQ coupler temperature increase. The system recovered and was reconditioned at low duty cycle with no issue. A maximum of 10% duty cycle was kept with varying pulse length and repetition rate. Upon request by INFN, field flatness measurement were performed at several duty cycle (with fix pulse duration of 4ms) before going to the next step increasing the duty cycle up to 15%. Until winter break the plan is to have beam operation in day time and keep conditioning during the night time and weekend.

Beam Transport/BD

Before resuming beam operation, there was a very intense period of remote sessions with CIEMAT to solve some last minute issue and complete few remaining tasks. Power Supply (PS) OPI was modified to allow the use of the PS HPU04 which is perfectly functional but does not provide digital current readback (analog output are operative and processed through the OPI). 

Diagnostics

In the remote sessions with CIEMAT the tasks for the cRIO SW update, the noise issues of BD accelerometer and the water temperature measurement at the D-slit were completed. BPM clock board was recovered by applying SW update.

Ancillaries/vacuum and cooling

Just before resuming the beam one vacuum gate valve (isolating a cryopump in the HEBT vacuum sector III) was blocked half open. Despite all the efforts it could not be recovered. A number of countermeasures have been taken to operate safely the vacuum system during the beam and the intervention on the valve for the next winter break is in preparation.

Improvements of the secondary cooling system (outdoor large water tank area) restarted to install protector against drop of ice and frozen snow.

The large Cryoplant 2nd storage buffer recently installed has been connected, conditioned and filled with some pure helium.

Control System

MPS test before beam operation was longer than planned and it was necessary to have CIEMAT to get all MPS signals for the DPlate diagnostic cleared. The MPS for the HEBT/BD was completed on 11 November in the morning.

Beam operation

A rehearsal for beam operation was done on 11 November and beam operation started on November 12.

Operations Supervisor and sub-systems operators observing traces of first beams in CCR
Remote participants observing traces of first beam on shared OPI

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (11/October-15/October)

Injector

The maintenance of injector in the vault was completed. Accelerator column alignment was checked and found to be in line with last year checks. Boron Nitride disks were found to be strongly damaged after half year operation: they have been replaced by new ones (after baking). The cause of damage (surface eroded and colored in black) needs to be identified. Spares Boron Nitride disks are available on-site.

The PPS test was performed on Tuesday and HV up to 100kV were applied with no issues. D+ plasma conditioning was resumed on Wednesday this week.

Radio Frequency (RF) System and RFQ cavity

The last activities before resume RF-RFQ modules were completed, namely the installation of the re-calibrated VSWR cards into the PSYS units and the installation of the Anode Boxes with new PCB circuits into RFQ module 1 and 4B. The start-up process of RFQ-RF modules started on Wednesday.

An accidental spill of water occurred from a 50 kW dummy load during a series of planned tests on the SSPA-A. To check the SSPA functionalities after the accident, tests procedures were prepared and tests carried out: the SSPA-A functionality is unaltered, with a nominal output power extracted from the SSPA modules. A damage was found in an electronic board of the peripheral control unit and spare procurement started.

On the RFQ manifolds in the vault, the check of water flow rates on the hundreds of individual channels equipped with flow regulators was completed and results confirmed that no immediate actions is needed.

Beam Transport

The checks and calibration of the HEBT magnet PS is almost completed.

The mechanical check of the MEBT bunchers has been completed.

Functional checks of mass flow controllers for HEBT vacuum pressure control is on-going with CIEMAT remote support.

Ancillaries

A new radiation resistant RGA and a new rupture disk were installed in vacuum sector IV of the HEBT.

All cryopumps for RFQ and HEBT have been successfully restarted.

Diagnostics

Newly calibrated HEBT FPMs (Fluorescence Profile Monitors) were installed. Functional check will be done though the remote session with CIEMAT.

Checks are ongoing on BD accelerometer cables with CIEMAT remote support.

Control system.

The integration of common chiller LCS into the CCS is on-going. The communication line between BD skid and HEBT LCS has been renewed.

In the current Integrated LIPAc schedule, it is planned to resume beam operation early November.

HEBT sector IV, new RGA radiation resistant + new rupture disk

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (27/Sepember-1/October)

As you may remember the LIPAc beam operation ended at the end of July.

Starting from August and until end of September there are major maintenance interventions in LIPAc, including regulatory inspections (on electrical system, cooling systems, HVAC, crane in the vault etc.) compulsory by law.

Injector

The new Emittance Meter Unit delivered by F4E has been installed in the vault, cleaned and the compatibility with the ancillaries and mechanical connections confirmed.

The main maintenance on the injector has just started and will continue until early October.

Injector conditioning is planned to resume by mid of October.

Radio Frequency (RF) System and RFQ cavity

The full maintenance of HVPS and RF-RFQ modules was performed with the support of NAT subcontractor. The RF team is currently working on upgrades of the RFQ RF modules to improve their functionality and stability in operation: upgraded PCB cards for the anode boxes have been validated and are being installed; PSYS VSWR cards to fit the new predrivers are under calibration; AC filters have been installed inside the PLC for noise reduction purposes.

By mid-October the RFQ RF modules will resume operation and target end of October to start injection in the RFQ. RF team will then need one week to perform three activities that needs RF injection in the cavity and then the RF will be ready for beam operation.

Beam Transport

Maintenance, checks and calibration of MEBT and MEL magnets is almost completed. HEBT magnets calibration is planned next week.

A new radiation resistance RGA and a new rupture disk will be installed this week in HEBT vacuum sector IV. The calibration and cable routing for the RGA are completed.

Ancillaries

A part the standard maintenance, several interventions were performed in the cooling skids (active carbons replacement and UPS replacement in the RF skid, valves inspections and recovery of functionality). Some pending interventions on the skids requires specific skills and a dedicated contract is being prepared by QST targeting the execution at the end of the year.

Also the newly commissioned Beam Dump skid underwent a thorough maintenance in September.

Diagnostics

Remote sessions with CIEMAT support have resumed to complete the commissioning of the HEBT and Beam Dump instrumentation (Faraday Cup, Fluorescence Profile Monitors, SEM grids motion control), debug HEBT PLC and complete the commissioning of the HEBT vacuum machine state. The pending hardware and cabling installation work was mostly executed onsite in August and remote sessions have been resumed in September

A VME control issues has to be solved to allow completing all the checks on Current Transformers and SEM grids data acquisition. CEA is supporting remotely.

A BPM in the MEBT is still not fully operational: the plan to remove the upper half of the quadrupole (MEBT BPM are embedded in the magnets) to access the BPM has been postponed to avoid the risk of the intervention.

Cryoplant

In September a large Helium buffer tank was installed using large cranes (see pictures below).

André Duarte was on-site for a month in July-August (first mission after a long time) to work on the integration in EPICS of the cryoplant (this included installation of a new PLC in the cubicles, connection to the local net and debugging).

Control system.

The Personal Protection system (PPS) annual maintenance and tests were carried out several weeks in September and are completed.

In the current Integrated LIPAc schedule, it is planned to resume beam operation early November.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (26/July-30/July)

Injector

Beam extraction with H+ at 50 keV, 20 mA for Phase B+ H+ commissioning was completed on 21 July 2021. D+ plasma conditioning was conducted from 22 to 25 July and D+ beam extraction at 100 keV, 30 mA was started on 26 July. D+ beam injection into RFQ started on 27 July.

Radio Frequency (RF) System and RFQ cavity

A short maintenance was performed for solving the issues on driver tetrode anode PS in chain 4B on 26 July (completed on 27 July morning). The RF system is operated stably at 1 Hz repetition (for 1 ms pulse). The RFQ conditioning towards 10% duty cycle has been put on-hold to prioritize beam operation.

Beam Instrumentation

A remote session with Ciemat was conducted on 26 July to improve the HEBT SEM Grid control. BPMs were not available at the beginning of the operations: they are progressively back on line and further checks are ongoing (it is necessary to debug a recent software update).

Control System

The Cryoplant LCS (Local Control System, stand-alone system at present) is under integration into LIPAc control system. We were happy to welcome after a long time an EU expert that will support this activity on-site in July and August.

LIPAc beam operation: H+ and D+ beam operation

The first H+ beam campaign at low current and low duty cycle has been successfully completed end of last week, fulfilling all the objectives defined in the Experimental Program, namely:

  • Injector test @20 mA extraction current (~10 mA at the exit of the LEBT)
  • Chopper test (100 msec)
  • LEBT (Low Energy Beam Transport) solenoid optimization (to maximize RFQ ACCT – AC Current Transformer- reading)
  • Transport up to HEBT FC (High Energy Beam Transport Faraday Cup)
  • Transport up to Beam Dump (monitor ACCTs) (~8 mA at the entrance)
  • Optimization of the dipole field
  • Beam dimension (by DPlate slits and HEBT SEM grid) and profile measurement (by HEBT SEM grids)
  • Alignment checks (screening with quadrupoles magnets)
  • Diagnostic checks

The D+ transport started beginning of this week. D+ extraction and transport up to the Beam Dump was successful with a good quality of the beam and excellent fraction of transported current (almost 80% from LEBT ACCT up to HEBT ACCT).

The following experimental program is being followed and target to be completed end of July:

  • Injector test @30 mA extraction current (~20 mA at the exit of the LEBT)
  • Chopper test (60 and 100 msec)
  • Source current adjustment
  • Transport up to HEBT FC
  • Transport up to Beam Dump (~16 mA at the entrance)
  • Optimization of the dipole field
  • Beam dimension (by DPlate slits and HEBT SEM grid) and profile measurement (by HEBT SEM grids)
  • Alignment checks (screening with quadrupoles magnets)
  • Buncher 2 phasing for BPM measurements and checks
  • Diagnostic checks as much as possible
  • Check of the beam barycenter and tuning of steerers

Next week the summer maintenance period will start. LIPAc operation is scheduled to resume end of September.

H+ beam transport along the LIPAc and beam current measurements along the line
D+ beam transport along the LIPAc and beam profile measurement at the SEM grid.
D+ beam transport along the LIPAc and beam current measurements along the line

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.